### v1

# Tensor rank and dimension expanders

Abstract
We prove a lower bound on the rank of tensors constructed from families of linear maps that ‘expand’ the dimension of every subspace. Such families, called dimension expanders have been studied for many years with several known explicit constructions. Using these constructions we show that one can construct an explicit $D]\times[n]\times[n]-$tensor with rank at least $(2-\epsilon)n$, with $D$ a constant depending on $\epsilon$. Our results extend to border rank over the real or complex numbers.
1 Overview
In the following, $\mathbb{F}$} denotes a field which is either finite or the real/complex numbers.
An $(n_{1},n_{2},n_{3})$-tensor over $\mathbb{F}$} is a function $T:[n_{1}]\times[n_{2}]\times[n_{3}]\rightarrow{\mathbb{F}}$. We can form two dimensional ‘slices’ of the tensor $T$ by fixing one of the variables. For example, we denote by T(a,,)T(a,*,*) the $n_{2}\times n_{3}$ matrix obtained by fixing $i_{1}$ to $a\in{\mathbb{F}}$. A tensor $T$ has rank one if it is of the form $T(i_{1},i_{2},i_{3})=f_{1}(i_{1})f_{2}(i_{2})f_{3}(i_{3})$ for some functions $f_{j}:[n_{j}]\rightarrow{\mathbb{F}}$. The rank of a general tensor $T$, denoted $R(T)$ is defined to be the minimal $r$ such that $T$ can be written as a sum of $r$ rank-one tensors. [^1] There are other notions of rank for tensors, all generalizing matrix rank, but the one we defined is typically called just ‘tensor rank’. When $\mathbb{F}$} is the real or complex numbers we also define the border rank of $T$, denoted $\underline{R}(T)$ to be the smallest $r$ such that there is a sequence of rank $\leq r$ tensors converging to TT. [^2] One can define border rank also over fields of positive characteristic as being in the Zariski closure of low rank tensors. While it is easy to show the existence of tensors of high rank (as high as $\Omega(n^{2})$ for $(n,n,n)$ tensors) it is a longstanding open problem to construct high-rank tensors explicitly. By explicit construction we mean that, there is a deterministic algorithm that, given $n$, runs in time $\textsf{poly}(n)$ and outputs an $(n,n,n)$-tensor $T_{n}$ of the required rank (more formally, an explicit construction produces an infinite family of tensors, not a single one). The difficulty in constructing explicit tensors of high rank stems from the fact that such constructions would lead to lower bounds on arithmetic circuits, a notoriously difficult open problem in theoretical computer science (see [SY10] for more details on this connection).
The currently best known explicit lower bounds on tensor rank are those in [AFT11] and are of the form $3n-O(\log_{2}(n))$ for an $(n,n,n)$-tensor (over any field $\mathbb{F}$}). Their work also includes a lower bound of 2n12n-1 on the rank of an explicit $(\log_{2}(n),n,n)$-tensor (for $n$ a power of two). For border rank the best explicit lower bound is that of [LM19] and is of the form $2.02n$ for an explicit $(n,n,n)$-tensor. To the best of our knowledge, the best tensor rank (or border rank) lower bounds for $(D,n,n)$-tensors with constant $D$ are of the form $3n/2$ [Gri78, Ja’78, Str83, Gri86]. In fact, one can get this bound even with $D=2$ (or $D=3$ for border rank).
Our main contribution is a new way to construct $(D,n,n)$-tensors of rank (or border rank) $(2-\epsilon)n$ with $D$ a constant only depending on $\epsilon$. We show a general way to construct such tensors from small families of linear maps that ‘expand’ or ‘spread’ the dimension of any subspace. The property we need is captured in the following definition.
Definition 1.1 (dimension-spreading).
Let ${\mathcal{L}}=\{L_{1},\ldots,L_{D}\}$ be a set of linear maps $L_{i}:{\mathbb{F}}^{n}\rightarrow{\mathbb{F}}^{n}$. We say that $\mathcal{L}$} is (s,t)-dimension-spreading if, for every subspace $U\subset{\mathbb{F}}^{n}$ of dimension $\dim(U)\geq s$ we have
$\dim\left(\sum_{i=1}^{D}L_{i}(U)\right)\geq t.$
We can extend the definition to sets of $n\times n$ matrices (thinking of them as linear maps).
To state our main theorem we need one additional notation. Given a set of $n\times n$ matrices ${\cal A}=\{A_{1},\ldots,A_{D}\}$ we define the $(D,n,n)$-tensor $T^{\cal A}$ to be
$T^{\cal A}(i,j,k)=(A_{i})_{jk}.$
In other words, $T^{\cal A}$ is the tensor composed of the $D$ ‘slices’ given by the matrices $A_{1},\ldots,A_{D}$.
Theorem 1.2.
Let $\mathbb{F}$} be any field. If ${\cal A}=\{A_{1},\ldots,A_{D}\}$ is a set of $n\times n$ matrices that is $(s,t)$-dimension spreading then the $(D,n,n)$-tensor $T^{\cal A}$ has rank at least n+tsn+t-s. Furthermore, if $\mathbb{F}$} is the real or complex numbers, then the same bound holds for border-rank as well.
We prove this theorem in Section 2. An interesting special case of this theorem is that of $(\epsilon n,(1-\epsilon)n)$-dimension spreading maps with $\epsilon$ a small constant. In this case, we can take $D$ to depend only on $\epsilon$ and the rank bound becomes $(2-2\epsilon)n$. We discuss explicit constructions of such dimension-spreading maps in Section 3. We observe that these can be constructed in a black-box manner from dimension-expanders for which there are several known explicit constructions [DS11, DW10, LZ08, BY13, FG15, GRX18]. In particular, using the construction in [BY13] one can get, for any $\epsilon$ and any sufficiently large $n$ an explicit family of $(\epsilon n,(1-\epsilon)n)$-dimension spreading maps over ${\mathbb{F}}^{n}$ of size $D=\textsf{poly}(1/\epsilon)$. [^3] We use the notation $\textsf{poly}(f)=O(f^{C})$ with $C$ some absolute constant. A nice property of this construction is that it is field independent – the same construction works over any field and the defining matrices have only zero and one entries. Hence we get the following corollary.
Corollary 1.3.
For every $\epsilon>0$, there is an explicit family of $(\textsf{poly}(1/\epsilon),n,n)$-tensors $T^{n,\epsilon}$ (defined using zeros and ones only) that has rank at least $(2-\epsilon)n$ over any field (or border-rank if $\mathbb{F}$} is the real or complex numbers).
2 Proof of Theorem 1.2
We start with a simple claim which restates the condition of having low tensor rank.
Claim 2.1.
Let $T$ be a $(D,n,n)$-tensor over $\mathbb{F}$} with $R(T)\leq r$. For each $i\in[D]$ let $A_{i}=T(i,*,*)$ be the $n\times n$ matrix obtained as a slice of $T$. Then, there exist $r$ rank-one $n\times n$ matrices $E_{1},\ldots,E_{r}$ that simultaneously span all the $A_{i}$’s. That is,
$\{A_{1},\ldots,A_{D}\}\subset\textsf{span}\{E_{1},\ldots,E_{r}\}.$
Proof.
Suppose $T=T^{1}+\ldots+T^{r}$ such that
$T^{\ell}(i,j,k)=f^{\ell}(i)g^{\ell}(j)h^{\ell}(k)$
is a rank-one tensor for all $\ell\in[r]$. We define the rank one matrices $E_{\ell}$ as follows:
$(E_{\ell})_{jk}=g^{\ell}(j)h^{\ell}(k).$
Fix some $i\in[D]$ and notice that
$(A_{i})_{jk}=T(i,j,k)=\sum_{\ell}T^{\ell}(i,j,k)=\sum_{\ell}f^{\ell}(i)g^{\ell}(j)h^{\ell}(k)=\sum_{\ell}f^{\ell}(i)\left(E_{\ell}\right)_{jk}.$
And so, $A_{i}$ is in the span of $E_{1},\ldots,E_{\ell}$ as was required. [^4] In fact, the proof shows that the two conditions are equivalent, but we will not use it.
To prove Theorem 1.2, suppose in contradiction that the rank of $T^{\cal A}$ was
$R(T^{\cal A})=r<n+t-s$ (1)
Let $E_{1},\ldots,E_{r}$ be the rank-one matrices given by Claim 2.1 spanning the matrices $A_{1},\ldots,A_{D}$.
For every subset $S\subset[r]$ we denote by
$K_{S}=\bigcap_{\ell\in S}{\mathbf{Ker}}(E_{\ell})$
and by
$I_{S}=\sum_{\ell\in S}{\mathbf{Im}}(E_{\ell}).$
Notice that for every set $S$ we have
$\dim(K_{S})\geq n-|S|,\,\,\,\,\,\dim(I_{S})\leq|S|.$ (2)
Claim 2.2.
For any set $S\subset[r]$ and for every map $A_{i}$, $i\in[D]$ we have
$A_{i}(K_{S})\subset I_{\overline{S}}.$
Proof.
Since
$A_{i}=\alpha_{1}E_{1}+\ldots+\alpha_{r}E_{r},$
for some coefficients $\alpha_{1},\ldots,\alpha_{r}\in{\mathbb{F}}$ we have that, if $v\in K_{S}$, then
$A_{i}(v)=\sum_{\ell\in\overline{S}}\alpha_{\ell}E_{\ell}(v),$
which is in $I_{\overline{S}}$ by definition. ∎
Let $S=\{1,2,\ldots,n-s\}$. Then
$|\bar{S}|=r-|S|<n+t-s-(n-s)=t.$ (3)
Now, from Claim 2.2, (2) and (3) we see that the $A_{i}$’s are not $(s,t)$-dimension-spreading, since
$A_{i}(K_{S})\subset I_{\bar{S}},$
$\dim(K_{S})\geq n-|S|=s$
and
$\dim(I_{\bar{S}})\leq|\bar{S}|<t.$
To argue about border rank, suppose that $\mathbb{F}$} is the real or complex numbers and suppose that there is a sequence of rank $\leq r$ tensors $T^{m}$ converging to $T^{\cal A}$. From the proof above, for each $m$ there will be some $s$ dimensional subspace $U^{m}$ such that
$\dim\left(\sum_{i=1}^{D}A_{i}^{m}(U^{m})\right)<t,$ (4)
Where $A^{m}_{1},\ldots,A^{m}_{D}$ are the $D$ slices of $T^{m}$ (so that $A^{m}_{i}$ converges to $A_{i}$ in the limit). By compactness of the Grassmannian, we can w.l.o.g assume that the sequence $U^{m}$ converges to some $s$ dimensional subspace $\hat{U}$. Now, we can pass to the limit in (4) to get that
$\dim\left(\sum_{i=1}^{D}A_{i}(\hat{U})\right)<t$
(using lower semi-continuity of matrix rank). Hence, we get that that $\hat{U}$ violates the $(s,t)$-dimension spreading condition for $\cal A$. This completes the proof of Theorem 1.2.
3 Constructing dimension-spreading maps
In this section we prove Corollary 1.3 by describing an explicit construction of an $(\epsilon n,(1-\epsilon)n)$-dimension spreading family of constant (depending on $\epsilon$) size. The construction will follow in a black box manner from the related (and well studied) notion of dimension expanders, defined by Barak, Impagliazzo, Shpilka and Wigderson [BISW04] as a linear algebraic analog of expander graphs.
Definition 3.1 (Dimension Expanders).
Let ${\mathcal{L}}=\{L_{1},\ldots,L_{D}\}$ be a set of linear maps $L_{i}:{\mathbb{F}}^{n}\rightarrow{\mathbb{F}}^{n}$. We say that $\mathcal{L}$} is a $\tau$-dimension-expander if, for every subspace $U\subset{\mathbb{F}}^{n}$ of dimension $\dim(U)\leq n/2$ we have
$\dim\left(\sum_{i=1}^{D}L_{i}(U)\right)\geq(1+\tau)\dim(U).$
We can extend the definition to sets of $n\times n$ matrices (thinking of them as linear maps).
As we will now show, it is quite straight-forward to construct an $(\epsilon n,(1-\epsilon)n)$-spreading family from an $\tau$-expanding one by taking sufficiently long words (compositions $L_{i_{1}}\circ L_{i_{2}}\circ$...) over the original family. One potential obstacle is that expansion is only guaranteed up to dimension $n/2$. We now show that this is not really an issue and expansion holds also for larger subspaces. For an $n\times n$ matrix $A$ we denote by $A^{*}$ its conjugate transpose (or just the transpose over finite fields).
Lemma 3.2.
Let ${\cal A}=\{A_{1},\ldots,A_{D}\}$ be a set of $n\times n$ matrices that is a $\tau$-dimension expander and such that $\cal A$ is closed under taking conjugate transpose and contains the identity. Let $U\subset{\mathbb{F}}^{n}$ be a subspace of dimension $\alpha n$ with $1/2<\alpha<1$. Then
$\dim\left(\sum_{i=1}^{D}A_{i}(U)\right)\geq\left(1+\frac{\tau(1-\alpha)}{2}\right)\dim(U).$
Proof.
Let $U$ be a subspace of dimension $\alpha n$ with $1/2<\alpha<1$. Let
$V=\sum_{i=1}^{D}A_{i}(U).$
Since we know that $\cal A$ contains the identity, there exists a non negative $\delta\geq 0$ such that
$\dim(V)=(1+\delta)\dim(U)=(1+\delta)\alpha n.$ (5)
Suppose $w\in V^{*}$, meaning
$\langle A_{i}^{*}w,u\rangle=\langle w,A_{i}u\rangle=0$
for all $u\in U$ and for all $i\in[D]$. Since $\cal A$ is closed under transpose conjugate, this implies that, if $w\in V^{*}$ then $\langle A_{i}w,u\rangle=0$ for all $u\in U$ and $i\in[D]$. This means that
$\sum_{i=1}^{D}A_{i}(V^{*})\subset U^{*}$
And so
$\dim\left(\sum_{i=1}^{D}A_{i}(V^{*})\right)\leq\dim(U^{*})=(1-\alpha)n.$ (6)
Combining (5) and (6) and using the dimension expansion property on $V^{*}$ (which has dimension at most $n/2$) we get that
$(1-\alpha)n\geq\dim\left(\sum_{i=1}^{D}A_{i}(V^{*})\right)\geq(1+\tau)\dim(V^{*})=(1+\tau)(1-(1+\delta)\alpha)n.$
solving for $\delta$ we get that
$\delta\geq\frac{\tau(1-\alpha)}{(1+\tau)\alpha}\geq\frac{\tau(1-\alpha)}{2}.$
We are now ready to show that sufficiently long words over a dimension expander gives a dimension spreading family. We will use the following notation: let ${\cal A}=\{A_{1},\ldots,A_{D}\}$ be a family of $n\times n$ matrices. For an integer $t\geq 1$ we denote by ${\cal A}^{(t)}$ the set of matrices obtained as a product of $t$ matrices from $\cal A$.
Lemma 3.3.
Let ${\cal A}=\{A_{1},\ldots,A_{D}\}$ be a set of $n\times n$ matrices that is a $\tau$-dimension expander and such that $\cal A$ is closed under taking conjugate transpose and contains the identity. Let $t>3\log_{2}(1/\epsilon)/\tau$. Then ${\cal A}^{(t)}$ is $(\epsilon n,(1-\epsilon)n)$-dimension spreading.
Proof.
Let $U=U_{0}$ be such that $\dim(U_{0})\geq\epsilon n$ and, for all $t\geq 1$ let
$U_{t}=\sum_{i\in[D]}A_{i}(U_{t-1}).$
Notice that as long as $\dim(U_{t})\leq n/2$ we have that
$\dim(U_{t+1})\geq(1+\tau)\dim(U_{t}).$
Thus, after at most $t_{1}\leq\tau^{-1}\log_{2}(1/\epsilon)$ steps we have $\dim(U_{t_{1}})>n/2.$ Once we are in this regime, we use Lemma 3.2 to argue about expansion. We can divide the interval $(1/2,1-\epsilon]$ into $\log_{2}(1/\epsilon)-1$ intervals
$J_{i}=(1-2^{-i},1-2^{-(i+1)}]$
for $i=1,\ldots,\log_{2}(1/\epsilon)-1$. By Lemma 3.2, if $\dim(U_{t})=\alpha n$ with $\alpha\in J_{i}$ then
$\dim(U_{t+1})\geq(1+\tau/2^{i+2})\alpha n.$
Hence, once $\alpha$ enters $J_{i}$, it will move to the next interval after at most $2/\tau$ steps. Indeed, suppose $\alpha>1-2^{-i}$ then
$\alpha(1+\tau/2^{i+2})^{2/\tau}>(1-1/2^{i})(1+1/2^{i+1})\geq 1-1/2^{i+1}.$
Hence, after an additional $2\log_{2}(1/\epsilon)/\tau$ steps we will have $\dim(U_{t})\geq(1-\epsilon)n$ as required. ∎
Hence, an explicit construction of an $\tau$-dimension-expander for some constant $\tau>0$ and with a constant number $D$ of matrices will imply an explicit construction of an $(\epsilon n,(1-\epsilon)n)$-dimension spreading family of size $D^{\prime}=\textsf{poly}(1/\epsilon)$. This, together with Theorem 1.2 (replacing $\epsilon$ with $\epsilon/2$) will imply Corollary 1.3. Notice that we can always include all the conjugate transposed (and the identity) with a constant blow up in $D$.
Constructions of Dimension Expanders:
There are three main approaches for constructing dimension expanders. The first, using representation theory, is due to Lubozky and Zelmanov [LZ08] and works over the complex numbers. What they show is that, if $G$ is a finite group generated as an expander by a set $S$ (i.e., the corresponding Cayley graph is a vertex expander) and if $\rho:G\mapsto U({\mathbb{C}}^{n})$ is an irreducible representation, then the set of linear maps $\rho(S)$ is a dimension expander (for some $\tau>0$ depending on the expansion of the associated Cayley graph). Another approach, yielding dimension expanders over sufficiently large finite fields, is given in [FG15, GRX18] and relies on ideas from coding/design theory.
Finally, the most general construction is due to Bourgain and Yehudayoff [BY13] and works over any field. It follows from an explicit construction of a constant degree monotone expander. A degree $D$ monotone graph is a bipartite graph $G$ on vertex set $n]\times[n]$ that can be partitioned into $D$ monotone matchings, which are (partial) matchings in which every pair of edges $(i_{1},j_{1})$ and $(i_{2},j_{2})$ satisfy $i_{1}<i_{2}\,\implies j_{1}<j_{2}$. If $G$ is also a vertex expander (every set $A\subset[n]$ of at most $n/2$ left vertices has at least $(1+\tau)|A|$ neighbors) then $G$ is a degree $D$ monotone expander. It was shown in [DW10] (and implicitly used in [DS11]) that a degree $D$ monotone expander gives a set of $D$ linear maps that are a $\tau$-dimension expander. The linear maps are defined directly from the matchings comprising the graph as follows: Suppose $f:[n]\rightarrow[n]\cup\{\bot\}$ is a partial function representing one of the monotone matchings such that $i_{1}<i_{2}\implies f(i_{1})<f(i_{2})$ (in case both are defined). We define the linear map $A_{f}$ sending $e_{i}$ to $e_{f(i)}$ if $f(i)\neq\bot$ and to 0 otherwise, where $e_{1},\ldots,e_{n}$ is the standard basis of ${\mathbb{F}}^{n}$. It is not hard to show that, over any field, the resulting family of $D$ maps $A_{f}$, taken over all matchings defining $G$, is a dimension-expander whenever the graph $G$ is a monotone expander.
Unlike with ‘standard’ expander graph (without monotonicity), there is no easy probabilistic argument showing existence of constant degree monotone expanders. In [DW10] explicit constructions of monotone expanders with slowly growing (but not constant) $D=D(n)$ were given using an analog of the Zig-Zag product. The construction in [BY13] is the only known construction (or even existence proof) of constant degree monotone expanders and uses a far reaching extension of the Bourgain-Gambourd ‘machine’ [BG07] (used to construct expanding Cayley graphs for finite groups) to the group $SL_{2}({\mathbb{R}})$. Using the dimension expanders of [BY13], which give some (unspecified but explicit) constants $D$ and $\tau$ and using Lemma 3.2 we conclude the proof of Corollary 1.3.

### v2

@@
 By explicit construction we mean that, there is a deterministic algorithm that, given $n$, runs in time $\textsf{poly}(n)$ and outputs an $(n,n,n)$-tensor $T_{n}$ of the required rank (more formally, an explicit construction produces an infinite family of tensors, not a single one).
-The difficulty in constructing explicit tensors of high rank stems from the fact that such constructions would lead to lower bounds on arithmetic circuits, a notoriously difficult open problem in theoretical computer science (see [SY10] for more details on this connection).
+The difficulty in constructing explicit tensors of high rank stems from the fact that such constructions would lead to lower bounds on arithmetic circuits, a notoriously difficult open problem in theoretical computer science (see [SY-survey] for more details on this connection).
 The currently best known explicit lower bounds on tensor rank are those in [AFT11] and are of the form $3n-O(\log_{2}(n))$ for an $(n,n,n)$-tensor (over any field $\mathbb{F}$}).
 Their work also includes a lower bound of 2n12n-1 on the rank of an explicit $(\log_{2}(n),n,n)$-tensor (for $n$ a power of two).
+It is not stated in [AFT11], but one can also slightly modify their Theorem 5.6 to construct a $(C,n,n)$-tensor with rank $(2-2^{-c})n$.
+As far as we can tell, the above mentioned results from [AFT11] do not extend to border rank.
 For border rank the best explicit lower bound is that of [LM19] and is of the form $2.02n$ for an explicit $(n,n,n)$-tensor.
-To the best of our knowledge, the best tensor rank (or border rank) lower bounds for $(D,n,n)$-tensors with constant $D$ are of the form $3n/2$ [Gri78, Ja’78, Str83, Gri86].
-In fact, one can get this bound even with $D=2$ (or $D=3$ for border rank).
+To the best of our knowledge, the best border rank lower bounds for $(D,n,n)$-tensors with constant $D$ are of the form $3n/2$ [Gri78, JaJa78, Str83, Grie86].
+In fact, one can get this bound even with $D=3$.
 Our main contribution is a new way to construct $(D,n,n)$-tensors of rank (or border rank) $(2-\epsilon)n$ with $D$ a constant only depending on $\epsilon$.
@@
 We can extend the definition to sets of $n\times n$ matrices (thinking of them as linear maps).
-To state our main theorem we need one additional notation. Given a set of $n\times n$ matrices ${\cal A}=\{A_{1},\ldots,A_{D}\}$ we define the $(D,n,n)$-tensor $T^{\cal A}$ to be
+To state our main theorem, we need one additional notation. Given a set of $n\times n$ matrices ${\cal A}=\{A_{1},\ldots,A_{D}\}$ we define the $(D,n,n)$-tensor $T^{\cal A}$ to be
 $T^{\cal A}(i,j,k)=(A_{i})_{jk}.$
@@
 We discuss explicit constructions of such dimension-spreading maps in Section 3.
-We observe that these can be constructed in a black-box manner from dimension-expanders for which there are several known explicit constructions [DS11, DW10, LZ08, BY13, FG15, GRX18].
+We observe that these can be constructed in a black-box manner from dimension-expanders for which there are several known explicit constructions [DS07, DW10, LZ08, BY13, FG15, GRX18].
 In particular, using the construction in [BY13] one can get, for any $\epsilon$ and any sufficiently large $n$ an explicit family of $(\epsilon n,(1-\epsilon)n)$-dimension spreading maps over ${\mathbb{F}}^{n}$ of size $D=\textsf{poly}(1/\epsilon)$.
@@
 A nice property of this construction is that it is field independent – the same construction works over any field and the defining matrices have only zero and one entries.
-Hence we get the following corollary.
+Hence, we get the following corollary.
 Corollary 1.3.
@@
 In this section we prove Corollary 1.3 by describing an explicit construction of an $(\epsilon n,(1-\epsilon)n)$-dimension spreading family of constant (depending on $\epsilon$) size.
-The construction will follow in a black box manner from the related (and well studied) notion of dimension expanders, defined by Barak, Impagliazzo, Shpilka and Wigderson [BISW04] as a linear algebraic analog of expander graphs.
+The construction will follow in a black box manner from the related (and well studied) notion of dimension expanders, defined by Barak, Impagliazzo, Shpilka and Wigderson [BISW] as a linear algebraic analog of expander graphs.
 Definition 3.1 (Dimension Expanders).
@@
 If $G$ is also a vertex expander (every set $A\subset[n]$ of at most $n/2$ left vertices has at least $(1+\tau)|A|$ neighbors) then $G$ is a degree $D$ monotone expander.
-It was shown in [DW10] (and implicitly used in [DS11]) that a degree $D$ monotone expander gives a set of $D$ linear maps that are a $\tau$-dimension expander.
+It was shown in [DW10] (and implicitly used in [DS07]) that a degree $D$ monotone expander gives a set of $D$ linear maps that are a $\tau$-dimension expander.
 The linear maps are defined directly from the matchings comprising the graph as follows: Suppose $f:[n]\rightarrow[n]\cup\{\bot\}$ is a partial function representing one of the monotone matchings such that $i_{1}<i_{2}\implies f(i_{1})<f(i_{2})$ (in case both are defined).
@@
 It is not hard to show that, over any field, the resulting family of $D$ maps $A_{f}$, taken over all matchings defining $G$, is a dimension-expander whenever the graph $G$ is a monotone expander.
-Unlike with ‘standard’ expander graph (without monotonicity), there is no easy probabilistic argument showing existence of constant degree monotone expanders.
+Unlike with ‘standard’ expander graph (without monotonicity), there is no easy probabilistic argument showing the existence of constant degree monotone expanders.
 In [DW10] explicit constructions of monotone expanders with slowly growing (but not constant) $D=D(n)$ were given using an analog of the Zig-Zag product.